# High-dimensional stochastic finite volumes using the tensor train format

### v1

@@ Abstract.
We propose a method for the uncertainty quantification of nonlinear hyperbolic equations with many uncertain parameters. The method combines the stochastic finite volume method and tensor trains in a novel way: the physical space and time dimensions are kept as full tensors, while all stochastic dimensions are compressed together into a tensor train. The resulting hybrid format has one tensor train for each spatial cell and each time step. We adapt a MUSCL scheme to this hybrid format and show the feasibility of the approach using several classical test cases. A convergence study is done on the Burgers’ equation with three stochastic parameters. We also solve the Burgers’ equation for an increasing number of stochastic dimensions and show an example with the Euler equations. The presented method opens new avenues for combining uncertainty quantification with well-known numerical schemes for conservation law.
@@ 1 Introduction
When adapting methods that were developed for low dimensions to high dimensions, the number of parameters at hand increases exponentially. This phenomenon is called the curse of dimensionality. Low-rank tensor formats appear as a possible solution to keep the storage and operations manageable. Low-rank tensors can be seen as a generalization to tensors of the well-known low-rank matrix decomposition. Several tensor formats have been developed, among them the canonical, the Tucker and the hierarchical formats [3]. Here, we focus on one hierarchical format, namely the tensor train format. Tensor trains have been successfully used for solving elliptic, parabolic and linear hyperbolic problems [4, 5, 6]. The non-linear hyperbolic case presents new difficulties because the structure is much less preserved over space and time: for example, shocks can appear. In [7], the authors propose to combine SFV with tensor trains and obtain promising results.
In this paper, we propose a mixed formulation where — in contrast to [7] — the time and physical space are kept in the full format while the stochastic space is compressed in the tensor train format. Section 2 introduces the problem and presents the principle of the stochastic finite volume method. In Section 3 we recall the main ideas of the tensor train format. The new mixed format is presented in Section 4. A MUSCL-type algorithm adapted to the mixed format is then described. It highlights the few modifications needed to use tensor trains compared to the classical scheme. In Section 5, we show numerical experiments that prove the feasibility and efficiency of our approach.
@@ 2.1 General description of the method
where ${\bf j}=(j_{1},\dots j_{m})$ is a multi-index. We introduce a constant mesh size in the spatial direction $\Delta x$, and in each stochastic dimension $\Delta\xi_{1},\dots,\Delta\xi_{m}$. The cells $K_{x}^{i}$ and $K_{\bm{\xi}}^{\bm{j}}$ are defined as
@@
and the cell average in space and in expectation over a cell ${\bf j}$
@@
$\int_{K_{\bm{\xi}}^{\bf j}}\int_{K_{x}^{i}}\frac{\partial{\bf u}}{\partial t}(t^{n},x,{\bm{\xi}})\;\text{d}x\,p({\bm{\xi}})\;\text{d}{\bm{\xi}}+\int_{K_{\bm{\xi}}^{\bf j}}\int_{K_{\bf x}^{i}}\frac{\partial{\bf f}}{\partial x}({\bf u}(t^{n},x,{\bm{\xi}}))\;\text{d}x\,p({\bm{\xi}})\;\text{d}{\bm{\xi}}={\bm{0}}.$ (9)
@@
[…]
The integrals are approximated using the midpoint rule. We introduce the approximated volume for the stochastic cells $K_{\bm{j}}=\Delta\xi_{1}\dots\Delta\xi_{m}\,p(\xi_{\bm{j}})$. The approximated average over a spatial cell is
$\bar{\bf u}_{i,{\bm{j}}}^{n}\approx u(t^{n},x_{i},{\bm{\xi}}_{\bm{j}})\,p({\bm{\xi}}).$ (12)
[…]
@@
[…]
@@ 2.2 Numerical flux
where intermediate values ${\bf u}_{i+1/2,{\bm{j}}}^{n+},{\bf u}_{i+1/2,{\bm{j}}}^{n-}$ are linear reconstructions of the solution using the approximate derivative, and $a_{i+1/2,{\bm{j}}}^{n}$ is the local speed. Since the midpoint rule is used to compute the expectation, all quantities should be understood as evaluated at the point ${\bm{\xi}}_{\bm{j}}$. The intermediate values are computed at the interface between two cells in space:
@@
The approximation of the derivative $({\bf u}_{x})_{i,{\bm{j}}}^{n}$ is computed with a limiter (here the minmod function),
@@ 2.3 Fully-discrete formulation and computational cost
$\left(\bar{\bf u}_{i,{\bm{j}}}^{n}\right)_{\begin{subarray}{l}n=1,\dots,N_{t}\\ i=1,\dots,N_{x}\\ {\bm{j}}=(1,\dots,1),\dots,(N_{\xi_{1}},\dots,N_{\xi_{j}})\end{subarray}}\in\mathbb{R}^{pN_{t}}\times\mathbb{R}^{pN_{x_{1}}}\dots\times\mathbb{R}^{pN_{x_{d}}}\times\mathbb{R}^{pN_{\xi_{1}}}\dots\times\mathbb{R}^{pN_{\xi_{m}}}.$ (21)
@@ 3.1 Definition of a tensor train
For $m\in\mathbb{N}$, an $m$th-order tensor $T~{}\in~{}\mathbb{R}^{N_{1}\times\dots\times N_{m}}$ is a tensor train if it can be written as a sum of products of third-order tensors,
$T(i_{1},\dots,i_{m})=\sum_{\alpha_{1}\dots\alpha_{m}=1}^{r_{1}\dots r_{m}}G_{1}(i_{1},\alpha_{1})G_{2}(\alpha_{1},i_{2},\alpha_{2})\cdots G_{m}(\alpha_{m},i_{m}).$ (22)
@@ 3.2 Operations on tensor trains
4 A mixed format for the stochastic finite volume method
@@
We propose here a method to solve hyperbolic equations with a high number of uncertain parameters. Our method takes advantage of the low-rank approximation while keeping the structure of deterministic solvers. The idea is to use the tensor train format only for the stochastic space, and to keep the full format in space and in time. We choose to keep a full format in time because we do not expect the solution to have a low-rank approximation with respect to time. The solution of a nonlinear hyperbolic problem can indeed significantly differ from the initial condition as time passes due to the steepening of gradients and, thus, discontinuities may develop. A comparison between low-rank approximation and full format in time was done in [6] for linear and quasi-linear hyperbolic problems. The authors observed that the full format in time was more accurate than its low-rank approximation, and we can expect the effect to be even more noticeable for non-linear problems. Moreover, keeping the full format in space and in time allows us to adapt more easily well-known deterministic solvers. Indeed, this way most steps are the same as a classical algorithm. The operations on scalars are simply replaced by operations on tensor trains, as we will describe in more detail later.
@@
The low-rank format is used in the stochastic directions, where the number of dimensions is the largest. There is no theoretical argument ensuring that a problem starting with a low-rank structure keeps a low-rank structure. In previous works [16, 17], it was even observed that shocks can appear in the stochastic direction, even if there is no flux in this direction. However, previous results [7] give experimental evidence that low-rank formats can well approximate the solution.
We now describe our method. For conciseness, the method is given for a scalar equation with one spatial dimension and with the same discretization in each stochastic dimension: $\Delta\xi_{1}=\Delta\xi_{2}=\dots=\Delta\xi_{m}=:\Delta\xi$, and $N_{\xi_{1}}=N_{\xi_{2}}=\dots=N_{\xi_{m}}=:N_{\xi}$. In this case, the discrete solution $\bar{u}$ has $N_{t}~{}N_{x}~{}N_{\xi}^{m}$ entries. We fix one time step $n$ and one spatial cell $i$, and consider the array of the numerical solution for all cells in the stochastic space. This $m$-dimensional array is approximated as a tensor train, denoted by $\bar{U}_{i}^{n}$.
$\bar{U}_{i}^{n}\approx\left(\bar{u}_{i,{\bm{j}}}^{n}\right)_{{\bm{j}}\in\{1,\dots,N_{\xi}\}^{m}}.$ (23)
By doing this for all spatial cells, we get a one-dimensional array (a vector) of tensor trains, one for each spatial cell. Hence, at the time step $n$ we consider the vector of tensor trains $\bar{U}^{n}=(\bar{U}_{1}^{n},\dots,\bar{U}_{N_{x}}^{n})^{\top}$. The same approximation is done for each reconstructed solution at the interface $i+1/2$. The resulting tensor train is denoted by $U_{i+1/2}^{n\pm}$.
$U_{i+1/2}^{n\pm}\approx\left(u_{i+1/2,{\bm{j}}}^{n\pm}\right)_{{\bm{j}}\in\{1,\dots N_{\xi}\}^{m}},$ (24)
and we consider the vector of tensor trains $U^{n\pm}=(U_{3/2}^{n\pm},\dots,U_{N_{x}+1/2}^{n\pm})^{\top}$.
With similar notations, let $\bar{H}^{n}=(\bar{H}^{n}_{1/2},\dots,\bar{H}^{n}_{N_{x}+1/2})$ be the vector of tensor trains representing the numerical flux expectation at the interfaces.
A practical example of our approach is now presented for the MUSCL scheme with forward Euler in time. Note that the idea is very general and can in principle be applied to any type of finite volume solver. After presenting the scheme, we describe how to compute the expectation and variance of the solution over the whole stochastic space.
4.1 The MUSCL scheme in the mixed format
@@
We adapt the scheme from Section 2 to apply it to arrays of tensor trains. Assume that, for a fixed cell $i$ and time step $n$, we know how to compute the tensor train approximations $\bar{U}^{n}_{i}$ and $U_{i+1/2}^{n+}$, $U_{i+1/2}^{n-}$. Then the scheme reads:
$\bar{U}_{i}^{n+1}$ $=\bar{U}_{i}^{n}-\frac{\Delta t}{\Delta x}t(\bar{H}_{i+1/2}^{n}-\bar{H}_{i-1/2}^{n}),$ (25)
@@
Equations (25)-(30) are similar to a classical MUSCL scheme with no stochastic variables. The only difference is that the objects $\bar{U}_{i}^{n},U_{i+1/2}^{n\pm},(U_{x})_{i}^{n}$ and $\bar{H}_{i+1/2}^{n}$ are scalars in the classical case, and tensor trains in our mixed formulation. Hence, all operations on these objects must be adapted to the tensor train format. As explained in Section 3, linear and polynomial functions can be applied directly to tensor trains, whereas the cross-interpolation algorithm must be used for non-polynomial functions. In our case, the operations $(T_{1},T_{2})\mapsto\text{minmod}(T_{1},T_{2})$ and the function $(T_{1},T_{2})\mapsto\max(|T_{1}|,|T_{2}|)$ are non-polynomial in the variables $T_{1},T_{2}$. Depending on the problem at hand, this can also be the case for the initial condition $u_{0}$, the flux $f$ and its derivative $f^{\prime}$.
The first step of the algorithm is to compute the cell average of the initial condition in the mixed format $\bar{U}^{0}$ using (12). For scalar equations, we make use of the maximum principle $|u(t,x,{\bm{\xi}})|\leq\max_{x,{\bm{\xi}}}|u_{0}(x,{\bm{\xi}})|$ to use the fixed time step
@@
The case for systems of equations will be described in Section 5.2. The solution is then propagated in time following the scheme (25)-(30).
The algorithms are summarized in Appendix A.

### v2

@@ Abstract.
A method for the uncertainty quantification of nonlinear hyperbolic equations with many uncertain parameters is presented. The method combines the stochastic finite volume method and tensor trains in a novel way: the physical space and time dimensions are kept as full tensors, while all stochastic dimensions are compressed together into a tensor train. The resulting hybrid format has one tensor train for each spatial cell and each time step. The MUSCL scheme is adapted to this hybrid format and the feasibility of the approach using several classical test cases is shown. For the Burgers’ equation a convergence study and a comparison with the full tensor train format are done with three stochastic parameters. The equation is then solved for an increasing number of stochastic dimensions. The Euler equations are then considered. A parameter study and a comparison with the full tensor train format are performed with the Sod problem. For a complex application we consider the Shu-Osher problem. The presented method opens new avenues for combining uncertainty quantification with well-known numerical schemes for conservation laws.
Keywords.
hyperbolic systems, stochastic finite volume method, tensor trains, low-rank approximation
MSCcodes.
65M08, 65M75, 35R60
@@ 1 Introduction
When adapting methods that were developed for low dimensions to high dimensions, the number of parameters at hand increases exponentially. This phenomenon is called the curse of dimensionality. Low-rank tensor formats appear as a possible solution to keep the storage and operations manageable. Low-rank tensors can be seen as a generalization to tensors of the well-known low-rank matrix decomposition. Several tensor formats have been developed, among them the canonical, the Tucker and the hierarchical formats [3]. Here, we focus on one hierarchical format, namely the tensor train format. Tensor trains have been successfully used for solving elliptic, parabolic and linear hyperbolic problems [4, 5, 6]. The nonlinear hyperbolic case presents new difficulties because the structure is much less preserved over space and time: for example, shocks can appear. In [7], the authors propose to combine SFV with tensor trains and obtain promising results.
In this paper, we propose a hybrid formulation where — in contrast to [7] — the time and physical space are kept in the full format while the stochastic space is compressed in the tensor train format. Section 2 introduces the problem and presents the principle of the stochastic finite volume method. In Section 3 we recall the main ideas of the tensor train format. The new hybrid format is presented in Section 4. A MUSCL-type algorithm adapted to the hybrid format is then described. It highlights the few modifications needed to use tensor trains compared to the classical scheme. In Section 5, we show numerical experiments that prove the feasibility and efficiency of our approach.
@@ 2.1 General description of the method
where ${\bm{j}}=(j_{1},\dots j_{m})$ is a multi-index. We introduce a constant mesh size in the spatial direction $\Delta x$, and in each stochastic dimension $\Delta\xi_{1},\dots,\Delta\xi_{m}$. The cells $K_{x}^{i}$ and $K_{\bm{\xi}}^{\bm{j}}$ are defined as
@@
and the cell average in space and in expectation over a cell $\bm{j}$}
@@
$\int_{K_{\bm{\xi}}^{\bm{j}}}\int_{K_{x}^{i}}\frac{\partial{\bf u}}{\partial t}(t^{n},x,{\bm{\xi}})\;\text{d}x\,p({\bm{\xi}})\;\text{d}{\bm{\xi}}+\int_{K_{\bm{\xi}}^{\bm{j}}}\int_{K_{\bf x}^{i}}\frac{\partial{\bf f}}{\partial x}({\bf u}(t^{n},x,{\bm{\xi}}))\;\text{d}x\,p({\bm{\xi}})\;\text{d}{\bm{\xi}}={\bm{0}}.$ (9)
@@
The exact flux $\bf f$} is replaced by a numerical flux $\mathbfcal F$: the value of the solution at the interface $x_{i+1/2}$ is approximated using the values of the averages in the surrounding cells $\bar{\bf u}_{i-1,{\bm{j}}}^{n},\bar{\bf u}_{i,{\bm{j}}}^{n},\bar{\bf u}_{i+1,{\bm{j}}}^{n}$. Let ${\mathbfcal F}_{{i},{{\bm{j}}}}^{{n}{+}}$ and ${\mathbfcal F}_{{i},{{\bm{j}}}}^{{n}{-}}$ denote the numerical flux at the interface $x_{i+1/2}$ and $x_{i-1/2}$, respectively. The approximation for their expectation over each cell is denoted by $\bar{\mathbfcal F}_{{i},{{\bm{j}}}}^{{n}{+}}$ and $\bar{\mathbfcal F}_{{i},{{\bm{j}}}}^{{n}{-}}$. Then Equation (10) is approximated by
$\Delta x|K_{\bm{\xi}}^{\bm{j}}|\frac{d\bar{\bf u}_{i,{\bm{j}}}^{n}}{dt}+\bar{\mathbfcal F}_{{i},{{\bm{j}}}}^{{n}{+}}-\bar{\mathbfcal F}_{{i},{{\bm{j}}}}^{{n}{-}}=0,\qquad\bar{\mathbfcal F}_{{i},{{\bm{j}}}}^{{n}{+}}=\mathbb{E}_{\bm{j}}[\mathcal{F}^{+}({\bf u}_{i,{\bm{j}}}^{n})],\quad\bar{\mathbfcal F}_{{i},{{\bm{j}}}}^{{n}{-}}=\mathbb{E}_{\bm{j}}[\mathcal{F}^{-}({\bf u}_{i,{\bm{j}}}^{n})].$ (11)
The integrals are approximated using the midpoint rule. The approximated average over a spatial cell is then
$\bar{\bf u}_{i,{\bm{j}}}^{n}\approx{\bf u}(t^{n},x_{i},{\bm{\xi}}_{\bm{j}})$ (12)
By abuse of notation, let $\bar{\bf u}_{i,{\bm{j}}}^{n}$ and $\bar{\mathbfcal F}_{{i},{{\bm{j}}}}^{{n}{\pm}}$ denote respectively the expectation of the solution and the numerical fluxes, both computed with quadrature. The midpoint rule is used to compute the expectation $\mathbb{E}$ with respect to $p$ and, therefore, quantities are evaluated at the point ${\bm{\xi}}_{\bm{j}}$. We introduce the approximated volume for the stochastic cells $K_{\bm{j}}=\Delta\xi_{1}\dots\Delta\xi_{m}\,p(\xi_{\bm{j}})$. The semi-discrete equation reads then
$K_{\bm{j}}\Delta x\frac{d\bar{\bf u}_{i,{\bm{j}}}^{n}}{dt}+K_{\bm{j}}\bar{\mathbfcal F}_{{i},{{\bm{j}}}}^{{n}{+}}-K_{\bm{j}}\bar{\mathbfcal F}_{{i},{{\bm{j}}}}^{{n}{-}}={\bm{0}}.$ (13)
@@
$\frac{d\bar{\bf u}_{i,{\bm{j}}}^{n}}{dt}+\frac{1}{\Delta x}\left(\bar{\mathbfcal F}_{{i},{{\bm{j}}}}^{{n}{+}}-\bar{\mathbfcal F}_{{i},{{\bm{j}}}}^{{n}{-}}\right)={\bm{0}}.$ (14)
@@ 2.2 Numerical flux
where intermediate values ${\bf u}_{i+1/2,{\bm{j}}}^{n+},{\bf u}_{i+1/2,{\bm{j}}}^{n-}$ are linear reconstructions of the solution using the approximate derivative, and $a_{i+1/2,{\bm{j}}}^{n}$ is the local speed.
The intermediate values are computed at the interface between two cells in space:
@@
The approximation of the derivative $({\bf u}_{x})_{i,{\bm{j}}}^{n}$ is obtained using a classical minmod limiter,
@@ 2.3 Fully-discrete formulation and computational cost
$\left(\bar{\bf u}_{i,{\bm{j}}}^{n}\right)_{\begin{subarray}{l}n=1,\dots,N_{t}\\ i=1,\dots,N_{x}\\ {\bm{j}}=(1,\dots,1),\dots,(N_{\xi_{1}},\dots,N_{\xi_{m}})\end{subarray}}\in\mathbb{R}^{pN_{t}}\times\mathbb{R}^{pN_{x_{1}}}\dots\times\mathbb{R}^{pN_{x_{d}}}\times\mathbb{R}^{pN_{\xi_{1}}}\dots\times\mathbb{R}^{pN_{\xi_{m}}}.$ (21)
@@ 3.1 Definition of a tensor train
For $m\in\mathbb{N}$, a $m$th-order tensor $T~\in~\mathbb{R}^{N_{1}\times\dots\times N_{m}}$ is a tensor train if it can be written as a sum of products of third-order tensors,
$T(i_{1},\dots,i_{m})=\sum_{\alpha_{1},\dots,\alpha_{m}=1}^{r_{1},\dots,r_{m}}G_{1}(i_{1},\alpha_{1})G_{2}(\alpha_{1},i_{2},\alpha_{2})\cdots G_{m}(\alpha_{m},i_{m}).$ (22)
@@ 3.2 Operations on tensor trains
4 A hybrid format for the stochastic finite volume method
@@
We propose here a method to solve hyperbolic equations with a high number of uncertain parameters. Our method takes advantage of the low-rank approximation while keeping the structure of deterministic solvers. The idea is to use the tensor train format only for the stochastic space, and to keep the full format in space and in time. We choose to keep a full format in time because we do not expect the solution to have a low-rank approximation with respect to time. The solution of a nonlinear hyperbolic problem can indeed significantly differ from the initial condition as time passes due to the steepening of gradients and, thus, discontinuities may develop. A comparison between low-rank approximation and full format in time was done in [6] for linear and quasi-linear hyperbolic problems. The authors observed that the full format in time was more accurate than its low-rank approximation, and we can expect the effect to be even more noticeable for nonlinear problems. Moreover, keeping the full format in space and in time allows us to adapt more easily well-known deterministic solvers. Most steps are indeed the same as a classical algorithm. The operations on scalars are simply replaced by operations on tensor trains, as we will describe in more detail later.
@@
The low-rank format is used in the stochastic directions, where the number of dimensions is the largest. There is no theoretical argument ensuring that a problem starting with a low-rank structure keeps a low-rank structure. In previous works [16, 17], discontinuities were even observed in the stochastic direction, although there is no flux in this direction. However, previous results [7] give experimental evidence that low-rank formats can well approximate the solution. The method introduced in [7] consists in compressing physical and stochastic spaces in a single tensor train. This single tensor train will be referred to as the "full-TT format" in the following.
Next, we discuss the proposed method in this manuscript. The method is stated for a scalar equation in a single spatial dimension and using the same discretization in each stochastic dimension. We define $\Delta\xi:=\Delta\xi_{1}=\Delta\xi_{2}=\dots=\Delta\xi_{m}$, and $N_{\xi}:=N_{\xi_{1}}=N_{\xi_{2}}=\dots=N_{\xi_{m}}$. A tensor train is used to approximate the solution for each spatial cell $i$, hence, the computational variable is an array of tensor trains:
$\bar{U}^{n}=(\bar{U}_{1}^{n},\dots,\bar{U}_{N_{x}}^{n})^{\top},\quad\text{ with }\quad\bar{U}_{i}^{n}\approx\left(\bar{u}_{i,{\bm{j}}}^{n}\right)_{{\bm{j}}\in\{1,\dots,N_{\xi}\}^{m}}.$ (23)
This hybrid-TT format is obtained by adding a new dimension to each core,
$\bar{U}^{n}_{i}({\bm{j}})=\sum_{\alpha_{1},\dots,\alpha_{m}=1}^{r_{1},\dots,r_{m}}G_{1}(i,j_{1},\alpha_{1})G_{2}(i,\alpha_{1},j_{2},\alpha_{2})\cdots G_{m}(i,\alpha_{m},j_{m}).$ (24)
For comparison we state the full-TT format. Therein, a single tensor train with a core for the spatial variable is used:
$\bar{U}^{n}(i,{\bm{j}})=\sum_{\alpha_{1},\dots,\alpha_{m}=1}^{r_{1},\dots,r_{m}}G_{0}(i,\alpha_{0})G_{1}(\alpha_{0},j_{1},\alpha_{1})G_{2}(\alpha_{1},j_{2},\alpha_{2})\cdots G_{m}(\alpha_{m},j_{m}).$ (25)
An example of the approach is presented for the MUSCL finite–volume scheme with a first–order explicit in time integration for the hybrid and full-TT scheme, respectively. However, the idea is general and may be applied to any finite-volume scheme. We also state the computation of the expectation of variance with respect to $p$ of the solution computed by the hybrid and full-TT schemes, respectively.
4.1 The MUSCL scheme in the hybrid format
@@
The scheme from Section 2 is adapted to the hybrid format. The reconstructed states and the expectation of the numerical flux in the hybrid format are respectively denoted by $U^{n\pm}=(U_{3/2}^{n\pm},\dots,U_{N_{x}+1/2}^{n\pm})^{\top}$, and $\bar{H}^{n}=(\bar{H}^{n}_{1/2},\dots,\bar{H}^{n}_{N_{x}+1/2})^{\top}$. We assume that, for a fixed cell $i$ and time step $n$, the tensor train approximations $\bar{U}^{n}_{i}$ and $U_{i+1/2}^{n+}$, $U_{i+1/2}^{n-}$ are given. Then, the MUSCL scheme in hybrid format reads:
$\bar{U}_{i}^{n+1}$ $=\bar{U}_{i}^{n}-\frac{\Delta t}{\Delta x}(\bar{H}_{i+1/2}^{n}-\bar{H}_{i-1/2}^{n}),$ (26)
@@
Equations (26)-(31) are similar to a classical MUSCL scheme with no stochastic variables. The only difference is that the objects $\bar{U}_{i}^{n},U_{i+1/2}^{n\pm},(U_{x})_{i}^{n}$ and $\bar{H}_{i+1/2}^{n}$ are scalars in the classical case, and tensor trains in our hybrid formulation. Hence, all operations on these objects must be adapted to the tensor train format. As explained in Section 3, linear and polynomial functions can be applied directly to tensor trains, whereas the cross-interpolation algorithm must be used for non-polynomial functions. In our case, the operations $(T_{1},T_{2})\mapsto\text{minmod}(T_{1},T_{2})$ and the function $(T_{1},T_{2})\mapsto\max(|T_{1}|,|T_{2}|)$ are non-polynomial in the variables $T_{1},T_{2}$. Depending on the problem at hand, this can also be the case for the initial condition $u_{0}$, the flux $f$ and its derivative $f^{\prime}$.
The first step of the algorithm is to compute the cell average of the initial condition in the hybrid format $\bar{U}^{0}$ using (12). The case for systems of equations will be described in Section 5.2. For scalar equations, we use the maximum principle $|u(t,x,{\bm{\xi}})|\leq\max_{x,{\bm{\xi}}}|u_{0}(x,{\bm{\xi}})|$ to set the fixed time step
@@
The solution is then propagated in time using the scheme (26)-(31). The algorithms are summarized in Appendix A.